# Learning local equivariant representations for quantum operators

### v3

@@ IV.2 Efficiency and Scalability
Figure 5: Comparison of training time per iteration and memory consumption with SLEM and DeepH-E3  [8] models.
Additionally, the strict local design assists the parallelization greatly. Since the edge and node features in the SLEM model are constructed locally, each node and edge feature is only dependent on the neighbour subgraph of the whole system, as:
$\mathbf{n}^{i,L}$ $=U_{L}\left(\mathbf{n}^{i,L-1},\sum_{j\in\mathcal{N}(i)}\mathbf{m}^{i,L}\right)$ (9)
$=U(Z_{i},\{\mathbf{Y}^{ij},Z_{j}\}_{j\in N(i)})$
Therefore, it is possible to divide the whole atomic system into sub-graphs and compute their node and edge features independently on different devices. This is extremely important when expanding the simulation power of DFT to very large systems. In practice, we find that for HfO2 with $4s2p2d2f1g$ basis, a typical model of 1.7M parameters can only predict the quantum operators of structures up to $10^{3}$ atoms on devices with 32GB memory. Despite the linear dependency, the inference on a system with $10^{4}\sim 10^{7}$ would require a memory with potentially over 300 GB, which is only available when parallelised to multiple GPUs. Therefore, a strictly localized model such as SLEM holds significant potential for expanding simulation system sizes.
Discussion
This work proposed SLEM, the strictly local equalvariant message-passing model for the prediction of quantum operator representations, as well as a systematic view of the parameterization of equivariant quantum operators. The strict local design of SLEM builds many-body dependence without expanding the effective receptive field, contributing to excellent accuracy, data efficiency and transferability. The integration of SO(2) convolution [19] enables rapid and memory-efficient tensor production, which is crucial for modeling heavy atoms. Furthermore, SLEM’s locality potentially allows for extension to extremely large atomic systems through parallelization over partitioned atomic graphs, thus enabling first-principles electronic structure simulations at device-level scales. Beyond its modeling capabilities, SLEM intrinsically supports the representation of multiple quantum operators, including the Hamiltonian, density matrix, and overlap matrix. By parameterizing the overlap matrix using a two-center integration formula, it is transformed into invariant quantities readily obtainable from SLEM’s initialized two-body embedding features. This approach significantly reduces computational costs and minimizes the model’s reliance on DFT software post-training.
  • Sampling Methodology for Active Learning: Scientific computation tasks require high confidence in calculated results, which can be challenging for data-driven approaches. Developing a robust sampling workflow for active learning is essential for reliability. While techniques like uncertainty-driven sampling with Gaussian regression [43] or model ensembles [44] have addressed confidence issues in machine learning force fields, designing a sampling workflow specifically for quantum operator models remains an open question.
  • Software Integration for Post-Processing: Integrating the model with existing software is vital, especially for high-throughput calculations and large atomic systems beyond conventional DFT capabilities. An efficient, parallelizable solver for extracting physical quantities from the model’s predictions is highly beneficial. While some software based on stochastic techniques  [45, 46] shows promise, an open-sourced and highly-optimized solution for non-orthogonal bases remains unavailable.

### v4

@@ 4.2 Efficiency and Scalability
Figure 5: Comparison of training time per iteration and memory consumption with SLEM and DeepH-E3  Gong et al. (2023) models.
In materials science, chemistry, and biology, many significant properties emerge in systems containing heavy atoms. These heavy atoms introduce high-order spherical tensors when representing quantum operators. Scaling to such systems is challenging due to the computational complexity of tensor products used to construct complex spherical tensors, which scales as $O(l^{6})$. Conventional tensor production methods struggle with training and inference on systems containing heavy atoms, making it difficult to model these important phenomena efficiently. Moreover, inferring large material systems while training with small structures is particularly valuable, which requires parallelising the model inference by assigning partitions of the large atomic structure to multiple GPU workers. However, most current models struggle with this task. As the receptive fields expand through iterative graph updates, the minimum size of each partitioned subgraph increases, reducing the effectiveness of such partitioning. The SLEM model addresses these challenges by efficiently constructing high-order tensor products and assisting parallelization through its strict locality.
For efficiency, the implementation of SO(2) convolution reduces the tensor product computational complexity from $O(l^{6})$ to $O(l^{3})$, which is then further reduced by the parallelization of matrix operations to nearly $O(l)$ benefiting from PyTorch. Figure 4 compares the wall time and GPU memory consumed by tensor product operations using SO(2) convolution versus the conventional method employed in DeepH-E3 Gong et al. (2023) and E3NN Geiger & Smidt (2022). The SO(2) convolution approach demonstrates markedly superior efficiency, enabling our method to handle all possible basis choices in LCAO DFT. We also evaluated memory usage and training time for typical systems, comparing our model with DeepH-E3. The results, displayed in Fig. 5, show that our model consistently outperforms in both metrics. Notably, the advantage becomes more pronounced as the basis set size increases, highlighting our method’s scalability. Additionally, the SLEM model’s strict locality design significantly enhances parallelization. This localized approach allows for the division of the atomic graph into sub-graphs, enabling independent computation of node and edge features on separate devices. This is extremely important when expanding DFT simulation to large systems. In practice, for HfO2 with $4s2p2d2f1g$ basis, a typical model of 1.7M parameters can predict the quantum operators for up to $10^{3}$ atoms on devices with 32GB memory. Despite linear scaling, the inference on a system with $10^{4}\sim 10^{7}$ would require over 300 GB of memory, necessitating parallelization across multiple GPUs. Therefore, a strictly localized model such as SLEM holds significant potential for expanding simulation system sizes.
5 Conclusion
This work presents SLEM (Strictly Local Equivariant Message-passing) model, a novel approach for predicting quantum operator representations in materials science. By employing a strict locality design and integrating SO(2) convolution, SLEM achieves state-of-the-art performance in predicting Hamiltonians, density matrices, and overlap matrices for diverse materials, including systems with heavy atoms. The model’s efficiency and scalability open new possibilities for large-scale quantum simulations and high-throughput materials discovery. Notably, SLEM’s locality enables efficient parallelization through atomic graph partitioning, potentially extending its applicability to extremely large systems at device-level scales. SLEM’s intrinsic support for multiple quantum operators, coupled with its novel overlap matrix parameterization, significantly reduces computational costs and dependence on post-training DFT software. The model demonstrates superior data efficiency and transferability, making it particularly well-suited for developing universal DFT models, especially for systems involving heavy elements. These advancements position SLEM as a powerful tool for simulating complex systems in materials science and computational chemistry. Looking ahead, future work will focus on developing robust sampling methodologies for active learning and integrating SLEM with existing software ecosystems to fully leverage its capabilities in real-world applications, further advancing the field of computational materials science.
@@ Acknowledgments
Appendix A Benchmark on QH-9
We also train our model on the open molecule Hamiltonian and Overlap dataset, QH-9 Yu et al. (2024), to test our method’s accuracy and transferability on a large dataset. We compare our result with the reported benchmark QH-net, which shows considerable accuracy improvement, as displayed in LABEL:qh9. We should notice that our model performs well in the Hamiltonian learning tasks, by decreasing 1/3 of the error reported previously. The density matrix is inferred from the learned Hamiltonian since the data is not available in QH9. Our accuracy is only 1/3 of the compared one.
QH9 is an extended electronic structure dataset based on QM9 Ruddigkeit et al. (2012); Ramakrishnan et al. (2014), a well-known benchmark for molecular property prediction. QH-9 contains 13,081 static structures and 2998 molecular trajectories. We test our method based on the static out-of-distribution(OOD) tasks defined by QHnet’s paper, where the training set and testing set are divided fully, where the testing set has a very different molecular structure and atom number w.r.t. training set. For OOD, the dataset is split as 104,001/17,495/9,335. As discussed in G, we used a semi-local variant of SLEM to prevent the possible break of strict locality in the molecular system.
Appendix B On the strict locality approximation
Table 6: MAE for Hamiltonian matrix (H) and density matrix (D) prediction.
QH-9 dataset benchmark
Method H[$10^{-6}E_{h}$] D
all non-diag diag all
LEM 56.57 44.86 155.24 0.0219
QHnet 83.22 80.26 135.63 0.0643
Here we want to further discuss the strict locality approximation made in our method. In the quantum operator learning task, we express the physical operators using an LCAO basis. In this expression, the theoretical framework of DFT fundamentally supports the strict locality approximation. In LCAO-based DFT calculations, the Hamiltonian matrix elements exhibit inherent locality. This locality manifests through two mechanisms: the spatial decay of LCAO basis functions, the electron neuralized conditions of the physics system, and electronic screening effects especially in condensed matter systems. These effects make the dependency of the quantum operator elements approximately local. For periodic systems, particularly metals and semiconductors, electronic screening effectively reduces long-range Coulomb interactions. The screening length varies by material type around 4 Å in semiconductors like silicon, and Germanium, and smaller in insulators such as 2.76 Å in Diamond Therefore, we are safe to choose a cutoff radius to match the sum of basis orbital radii, adequately captures the correct physical relations. Taking the Hamiltonian matrix elements as an example:
[…]
The locality is a combination of the LCAO basis, Hartree term, external potential, and xc functional. Therefore, we need to discuss the localization as a collective effect, which is contributed by the following properties:
The decaying behaviour of the LCAO basis.
For all LCAO bases (whether GTO, NAO, or Slater-like), the radial part all decays rapidly with distance. Therefore, the long-range term in $V_{H},V_{ext},V_{xc}$’s contribution in the integration (that decays as 1/r) would be decreased fastly (often exponentially such as in GTO or Slater-like orbital).
The system’s electronic neutralization condition.
The electronic neutralization condition requires that the long-term part of the Hartree term and the external term cancel each other. To derive this, we assign the electron density to the atomic center, by writing as:
$n(r)=\sum_{I}n_{I}(r-R_{I})$
Then the sum of $V_{H}(r),V_{ext}(r)$ would be:
[…]
When $|r-R_{I}|$ is large, we can approximate $|r-r^{\prime}|\approx|r-R_{I}|$, therefore:
[…]
The screening effect.
we refer to  Huckel & Debye (1923); Resta (1977); Ninno et al. (2006) the study of screening radius $R_{s}$, which describes the system’s electrostatic potential’s reaction to the vibration of charges.
In Resta (1977), the screening radius of typical insulator and semiconductor systems is reported as: Diamond: 2.76 a.u, Silicon: 4.28 a.u. and Germanium: 4.71 a.u. In Ninno et al. (2006), for some nanoparticles, the $R_{s}$ is reported as: Si191H148: 5.36 a.u. and Ge191H148: 5.61 a.u.
These effects motivate us to design SLEM, as a strict localized model that is suitable to learn the correct mapping from structures to quantum operators with the correct physical priors. Consititude is a more reliable, data-efficient method. Meanwhile, we also notice the existing limitation of this method in some special systems, we refer to the discussion in G for detail.
Appendix C Comparison of SLEM with MPNN
Message-passing Neural Networks
In the message-passing scheme, atoms are treated as nodes in graphs, with bonds to neighbouring atoms represented as connected edges within a specified cutoff radius. The embedded atomic features are processed by trainable functions, generating messages from each edge to update the embeddings of central atoms. Formally, the MPNN framework can be summarized as follows:
Here, $\mathbf{e}^{ij,L}$ represents the edge features, $\mathbf{m}^{ij,L}$ denotes the messages, and $\mathbf{n}^{i,L}$ indicates the node features at layer $L$. $\mathcal{N}_{L}$, $M_{L}$, and $U_{L}$ are the trainable functions for the edge, message, and node updates. $\mathcal{N}(i)=\{j|r_{ij}<r_{\text{cut}}\}$ indicates all the neighbour atoms for atom $i$, with $r_{\text{cut}}$ being the predefined cutoff. This updating framework allows for the construction of many-body interactions and long-term dependencies, leading to strong performance across various applications. Here we reframe the updating rules of SLEM with the MPNN framework, which looks like this:
$\mathbf{e}^{ij,L}$ $=\mathcal{N}_{L}\left(\mathbf{n}^{i,L-1},\mathbf{V}^{ij,L},\mathbf{n}^{j,L-1},\mathbf{e}^{ij,L-1}\right)$
$\mathbf{m}^{ij,L}$ $=M_{L}\left(\mathbf{n}^{i,L-1},\mathbf{V}^{ij,L}\right)$
Here $\mathcal{V}_{L}$ is the neural network for hidden feature construction. This update scheme constructs many-body interactions to build equivariant edge and node features while preserving the absolute locality by excluding atoms outside the constant cutoff radius. Such a scheme, in principle, would have much better transferability, data efficiency, as well as scalability when we have a strong prior that the input and output are dependent spatially local.
Appendix D More on tensor product
$\begin{split}\mathbf{f}^{ij}_{c_{3},l_{3}}&=\tilde{\mathbf{f}}^{ij}_{c_{1},l_{1}}\otimes w^{ij}_{c_{2},l_{2}}\mathbf{Y}_{l_{2}}^{ij}=\sum_{c_{1},l_{1},l_{2}}\tilde{w}_{c_{1},l_{1},l_{2}}^{ij}\sum_{m_{1},m_{2}}C_{(l_{1},m_{1})(l_{2},m_{2})}^{(l_{3},m_{3})}f^{ij}_{c_{1},l_{1},m_{1}}Y_{l_{2},m_{2}}^{ij}\end{split}$ (18)
Here, $\tilde{w}_{c_{1},l_{1},l_{2}}^{ij}=\sum_{c_{2}}w_{c_{1},c_{2},l_{1},l_{2}}w^{ij}_{c_{2},l_{2}}$ are edge-specific parameters for each tensor product operation. Performing such tensor products on high-order features is computationally intensive. Therefore, we applied the recently developed SO(2) Passaro & Zitnick (2023) convolution to simplify, reducing the computation and storage complexity from $O(l_{\text{max}}^{6})$ to $O(l_{\text{max}}^{3})$. The simplification idea is intuitive. $\mathbf{Y}_{l_{2},m_{2}}^{ij}$ are sparse tensors if rotated to align with the edge $ij$, which is nonzero only for $m_{2}=0$. Therefore, it is easier to compute the tensor production in the direction of edge $ij$, and rotate inversely the output afterwards. This step removes the $m_{2}$ index from the summation in Eq. 18. Furthermore, considering the Clebsch-Gordan coefficients with $m_{2}=0$, we find that $C_{(l_{1},m_{1})(l_{2},0)}^{(l_{3},m_{3})}=0$ except for $m_{3}=\pm m_{1}$. This allows further reduction of the summation in Eq. 18 by replacing ±$\pm m_{1}$ with a single index $m$. Then the operations can be reformulated formally as:
Appendix E Data Generation
In this section, we discuss the data generation process used in experiments of this work. The data sampling process includes the Hamiltonian, overlap and density matrix of materials Si, GaN, HfO2.
First, we perform an ab-initio molecular dynamic simulation using a neural network force field using DeePMDWang et al. (2018) and LammpsThompson et al. (2022). A typical input file of the lammps sampling looks like:
variable        NSTEPS          equal 500000
variable        THERMO_FREQ     equal 100
variable        DUMP_FREQ       equal 1000
variable        TEMP            equal 300
variable        PRES            equal 1.00
variable        TAU_T           equal 0.10
variable        TAU_P           equal 0.50
units           metal
boundary        p p p
atom_style      atomic
neighbor        1.0 bin
read_data       conf.lammps-data
mass            1 28.085
pair_style      deepmd graph.pb
pair_coeff      * * Si
thermo_style    custom step temp pe ke etotal press vol lx ly lz xy xz yz
thermo          \${THERMO_FREQ}
restart         1000000 dpgen.restart
velocity        all create \${TEMP} 449414
fix             1 all nvt temp \${TEMP} \${TEMP} 0.1
dump            1 all custom \${DUMP_FREQ}  eq.lammpstrj id type x y z vx vy vz
timestep        0.001
run     100000
undump 1
dump            2 all custom \${DUMP_FREQ}  sample.lammpstrj id type x y z vx vy vz
run             \${NSTEPS} upto
The samples are taken after $10^{5}$ steps when the system reaches equilibrium.
After the configuration sampling, we then applied ABACUSLi et al. (2016) to compute the quantum operators, including the Hamiltonian, density and overlap matrix of each configuration. Among them, 150 frames are randomly split as the training set, and 30 frames are used for testing. We use SG15 ONCV numerical atomic orbitals, pseudopotentialsHamann (2013), and PBE functionalsErnzerhof & Scuseria (1999) for the DFT calculation. A typical DFT calculation input looks like:
INPUT_PARAMETERS
# Created by Atomic Simulation Enviroment
ntype                                   1
ecutwfc                                 100
scf_nmax                                100
smearing_method                         gaussian
smearing_sigma                          0.002
basis_type                              lcao
ks_solver                               genelpa
mixing_type                             pulay
mixing_beta                             0.7
scf_thr                                 1e-07
out_chg                                 1
symmetry                                1
calculation                             scf
out_band                                1
force_thr                               0.001
out_stru                                1
kspacing                                0.08
lspinorb                                0
out_wfc_lcao                            0
dft_functional                          pbe
out_mat_hs2                             True
After the calculation converged for all, we processed the data format into machine-learning readable datasets and started training. The dataset used in this work is uploaded in the opensource platform AISquare via this link: https://www.aissquare.com/datasets/detail?pageType=datasets&name=Quantum_Operator_Dataset&id=286.
Appendix F Cutoff Selection
The cutoff used in the SLEM model is important since it decides the size of the atom and bond environment to consider when learning the map from the atomic structure to the quantum operators. There are two cutoff concepts in the SLEM model. One is the cutoff for the bond $r_{b}$, which is used to build the one-to-one correspondence of the graph structure of atomic data to the diagonal and off-diagonal blocks of the quantum operators. Another one is the cutoff for environment $r_{e}$.
For $r_{b}$, since the bond whose radial distance beyond $r_{b}$ would all be considered 0, this value should be fixed according to the LCAO orbital used in DFT data generation. For numerical atomic orbitals (NAOs), the cutoff is defined clearly in orbital files or DFT software, therefore, we just double (since two orbitals constitute a bond) and assign the value to each atom as there $r_{b}$. For Gaussian-type orbitals (GTOs) and Slater-type orbitals (STOs), since they decay exponentially along radial distance, we often set some precision threshold during data processing and find the largest bond in off-diagonal blocks that is larger than the threshold value. The bond length will be considered as our $r_{b}$.
The environmental cutoff $r_{e}$ is very system dependent. This decides how large atomic/bond environments relate to the mapping. This should follow the locality behaviour of the physical system. For example, in typical periodic crystals such as semiconductors (Si, GaN, etc.), medals (Graphene), and insulators (HfO2) the $r_{e}$ are quite local, therefore, it is safe to set the $r_{e}=r_{b}$. However, when systems experience strong nonlocality behaviour, $r_{e}$ should be very large. One should carefully balance the efficiency of choosing a large $r_{e}$ to ensure strict locality and the usage of a semi-local or MPNN-based method.
Appendix G Limitation
The SLEM model, benefiting from the Strictly localized design, performs better in accuracy and transferability as demonstrated in various datasets. We acknowledge that such a localized hypothesis is most suitable for describing periodic systems, which is commonly adopted in physics, chemistry and material science research. For confined systems such as molecules, the absence of screening effect could lead to long-term dependency that is uncovered within prefixed cutoff. In these cases, the SLEM model wouldn’t perform as well as in periodic cases. For generality, we also proposed a semilocal model called LEM (Localized Equivarient Message-Passing), where the interaction between distant atoms is included but decays exponentially with their distance, with a trainable decay factor. All the designed models are ready to use in our GitHub repository.
Appendix H Future Investigation
  • Sampling Methodology for Active Learning: Scientific computation tasks require high confidence in calculated results, which can be challenging for data-driven approaches. Developing a robust sampling workflow for active learning is essential for reliability. While techniques like uncertainty-driven sampling with Gaussian regression Vandermause et al. (2020) or model ensembles Zhang et al. (2020) have addressed confidence issues in machine learning force fields, designing a sampling workflow specifically for quantum operator models remains an open question.
  • Software Integration for Post-Processing: Integrating the model with existing software is vital, especially for high-throughput calculations and large atomic systems beyond conventional DFT capabilities. An efficient, parallelizable solver for extracting physical quantities from the model’s predictions is highly beneficial. While some software based on stochastic techniques  Li et al. (2023); João et al. (2020) shows promise, an open-sourced and highly-optimized solution for non-orthogonal bases remains unavailable.